OpenStudy (anonymous):

solve the equation? (2x+7) (7x-2)=0

OpenStudy (anonymous):

2x + 7 = 0, 2x = -7 therefore x = -7/2 7x-2= 0 7x = 2 therefore x=2/7

OpenStudy (anonymous):

x=0.286

OpenStudy (anonymous):

so x = 2/7 or x= -7/2

OpenStudy (anonymous):

the winner is rahulp10 thank u for the help

OpenStudy (anonymous):

You're welcome :)